OpenStudy (anonymous):

The first five terms of a sequence are shown below. 4, 7, 10, 13, 16 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence? A. f(n) = 3n - 1; n ≥ 1 B. f(n) = 3n + 1; n ≥ 1 C. f(n) = 4n - 1; n ≥ 1 D. f(n) = 4n + 1; n ≥ 1

OpenStudy (anonymous):

@uri @triciaal @aum @MakaylaTracy

OpenStudy (anonymous):

a+d(n-1) Is the formula A= the first term D= what it's going up by substitute in numbers and simplify

OpenStudy (cwrw238):

try plugging in 1 ,2 ,3 for n into each of the options

OpenStudy (michele_laino):

please note that your function is an arithmetic progression with first term = 4 and constant=3 so n-term is: \[f(n)=a _{n}=a _{1}+(n-1)*3=4+3n-3=3n+1\]

OpenStudy (anonymous):

@pooja195 @SolomonZelman @jacobwat @dtuniyants

OpenStudy (anonymous):

@AnswerMyQuestions @AnswerMyQuestionsPlease @aum